OpenStudy (anonymous):

can someone look at this problem and explain what to do with this last step, i am lost...thanks it is all posted..

OpenStudy (anonymous):

\[\frac{ x ^{2}+9x+20 }{x ^{2}-25 ? }\div \frac{ x+4 }{x-4 ? }\]

OpenStudy (anonymous):

I know to change division to multiply and flip \[\frac{ (x+4)(x+5 }{x ^{2}-25 }\times \frac{ x-4 }{x+4 ? }\]

OpenStudy (anonymous):

got to here, now my problem, what do I do with the \[x ^{2}-25\]

ganeshie8 (ganeshie8):

hint : \(25 = 5^2\)

ganeshie8 (ganeshie8):

hint2 : \(a^2-b^2 = (a+b)(a-b)\)

OpenStudy (anonymous):

(x+5)(x-5)

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

ok, now i cross multiply and get rid of (x+4) and the (x+5)...leaving \[\frac{ 1 }{ x-5? }*x-4\]

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

so the answer is \[\frac{ x-4 }{x-5 }\] correct?

ganeshie8 (ganeshie8):

yes ! good work !!

OpenStudy (anonymous):

thanks, i like doing them myself, just need a little clarification sometimes ;)

ganeshie8 (ganeshie8):

thats nice =)